OpenStudy (anonymous):

the length of a rectangle is 2 inches longer than the width. the area of the rectangle is 56. find the dimensions

OpenStudy (wolf1728):

L * W = 56 L =W+2 L -W -2=0 From first equation L = 56 / W 56/W -W -2 = 0 56 -W² -2W =0 -W² -2W +56 = 0 W² + 2W -56 =0 W=-2 +- sqrt(4 -4*1*-56) / 2*1 W = -2 +-sqrt(228)/2 W = 6.5498 L = 8.5498